OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6.

OpenStudy (anonymous):

x=7/4 or 1.4

OpenStudy (petiteme):

Use distributive property :) 3x+3 = -2x +2 +6 3x+3 = -2x+8 5x=5 divide both sides by 5 x=1

OpenStudy (anonymous):

shes right

OpenStudy (anonymous):

x=1